# AI-Generated Video Detection via Perceptual Straightening

### v1

@@ 10.3 Qualitative Examples of “Plants” Task
10.4 Results Distributions for Main Task (M) and Plants Task (P)
@@
11 Ablation Studies
@@
11.1 Impact of Video Length and Sampling Density
@@
11.2 Robustness to Temporal Window Position
@@
12 Computational Environment
@@
13 Dataset Licenses and Sources
  • VidProM [3]: This dataset was employed for training our video classifier (Section 6) and for benchmarking in Section 7.1 and Section 7. The VidProM dataset is offered for non-commercial research purposes under CC BY-NC 4.0 license.
  • GenVidBench [17]: GenVidBench was used for benchmarking in Section 7). The GenVidBench dataset and its associated code are under CC BY-NC-SA 4.0 license.
@@ 13 Dataset Licenses and Sources
  • DVSC2023 (Natural Video Source for Classifier Training): As explained in Section 5, a set of 50,000 natural videos for training (Section 6) was sourced from DVSC2023 [45]. DVSC2023 is under the CC BY-SA 4.0 llicense.

### v2

@@ A.4 Qualitative Examples of “Plants” Task
A.5 Results Distributions for Main Task (M) and Plants Task (P)
@@
A.6 Frame Samples from the Veo3 Model for Zero-shot Generalization Test
Figure 14: Qualitative frame examples from the Veo3 [66] model. These images showcase the high-fidelity content used for the zero-shot generalization test.
To provide a qualitative sense of the videos used in our zero-shot generalization test (Section 7, Paragraph D), Figure˜14 presents sample frames generated by the Veo3 model. These examples illustrate the high quality of the content our method was tested against without any prior training on this specific generator.
Appendix B Ablation Studies
@@
B.1 Impact of Video Length and Sampling Density
@@
B.2 Robustness to Temporal Window Position
@@
Appendix C Analysis of Scene Cut Frequency
Model Avg. Duration (s) Avg. Cuts in first 2s
Pika 2.97 0.00
Show-1 3.62 0.00
ModelScope 3.84 0.00
Gen-2 4.00 0.00
HotShot 4.60 0.01
Crafter 5.43 0.00
MoonValley 5.59 0.00
T2VZ 6.01 0.00
VC2 6.98 0.10
Natural 7.04 0.08
MorphStudio 7.05 0.02
OpenSora 8.22 0.04
Sora 12.40 0.04
Figure 17: Analysis of Scene Cut Frequency. The lollipop plot (left) and table (right) show the average number of hard scene cuts within the first 2s for each video source. The dashed blue line in the plot indicates the rate for natural videos. Both visualizations confirm that the scene cut frequency is low across all models and not a significant confounding factor.
An important consideration for AI-generated video detection is the effect of hard scene cuts, as the straightening hypothesis is not expected to hold across shot boundaries. This presents a potential confound: if AI-generated videos simply contained a higher frequency of scene cuts, it could partly explain our classifier’s performance.
To investigate this possibility, we analyzed 13,000 AI-generated and 13,000 natural videos using the “scenedetect” [107] library. The results, presented in Fig.˜17, show that the average number of scene cuts is low and comparable for both natural and AI-generated videos. The plot (left) visualizes this comparison, showing all models clustering near the natural video baseline (dashed line), while the table (right) provides the precise data.
Our method’s reliance on a short, 2s analysis window inherently reduces the probability of encountering a scene cut. The overall robustness of our approach is further confirmed by the ablation study in Section˜B.2, which shows stable detection performance regardless of the analysis window’s temporal position.
Appendix D Computational Environment
@@
Appendix E Dataset Licenses and Sources
  • VidProM [3]: This dataset was employed for training our video classifier (Section˜6) and for benchmarking in Section˜7 and Section˜7. The VidProM dataset is offered for non-commercial research purposes under $C$ BY-NC 4.0 license.
  • GenVidBench [17]: GenVidBench was used for benchmarking in Section˜7). The GenVidBench dataset and its associated code are under $C$ BY-NC-SA 4.0 license.
@@ Appendix E Dataset Licenses and Sources
  • DVSC2023 (Natural Video Source for Classifier Training): As explained in Section˜5, a set of 50,000 natural videos for training (Section˜6) was sourced from DVSC2023 [48]. DVSC2023 is under the $C$ BY-SA 4.0 llicense.